# ScrewSplat: An End-to-End Method for Articulated Object Recognition

### v1

@@ B.1 Details for Screw Theory
[…]
@@
where $T^{\prime}\in\mathrm{SE}(3)$ denotes the transformed rigid body coordinate, and $[\mathcal{S}]$ is the $\times 4$ 4 matrix representation of the screw axis $\mathcal{S}$, defined as
$[\mathcal{S}]=\begin{bmatrix}[\omega]&v\\ 0&0\end{bmatrix}\in\mathbb{R}^{4\times 4},\>\>\>[\omega]=\begin{bmatrix}0&-\omega_{3}&\omega_{2}\\ \omega_{3}&0&-\omega_{1}\\ -\omega_{2}&\omega_{1}&0\end{bmatrix}\in\mathbb{R}^{3\times 3},$ (12)
@@
For a revolute joint, $G(\theta)$ has a closed-form expression using the fact that $[\omega]^{3}=[\omega]$:
@@ B.2 Implementation Details for ScrewSplat
Initialization of Core Components. For the part-aware Gaussian primitives, we initially spawn 10,000 primitives. For each primitive $\mathcal{H}_{i}=(T_{i},s_{i},\sigma_{i},c_{i},m_{i})$, we generally adopt the same initialization scheme as used in the original Gaussian Splatting method for the variables $(T_{i},s_{i},\sigma_{i},c_{i})$. It is worth noting that the position $\mu_{i}$ of the Gaussian primitive from the pose $T_{i}=[R_{i},\mu_{i}]$ is sampled from a uniform distribution: in the simulation environment, it is sampled from $[-1,1]^{3}$, , and in real-world experiments, it is sampled from $[-0.7,0.7]^{3}$. We found that initializing the positions within the camera’s range results in better performance. The part probability $m_{i}\in\Delta^{n_{s}}$ is initialized as a uniform distribution (i.e., an element-wise constant vector).
For the screw primitives, we initially spawn eight revolute joint axes and eight prismatic joint axes (i.e., a total of 16 joint axes). For each screw primitive $\mathcal{A}_{j}=(\mathcal{S}_{j},\gamma_{j})$, we first sample a six-dimensional real vector from a uniform distribution over $[-0.5,0.5]^{3}$ and then normalize it to satisfy the constraints for both revolute and prismatic joints. Specifically, let the sampled six-dimensional vector be $[x,q]$, where $x\in\mathbb{R}^{3}$ and $q\in\mathbb{R}^{3}$. For a revolute joint $[\omega,v]$, $\omega$ is set to $x/||x||$ and $v$ is set to $-\omega\times q$. For a prismatic joint $[\omega,v]$, $\omega$ is set to the zero vector, and $v$ is set to $q/||q||$. The screw confidence $\gamma_{j}$ is initialized to 0.9.
@@ B.3 Details for Articulated Object Manipulation
[…]
@@
[…]
@@ C.1 Additional Details for Recognition Experiments
[…]

### v2

@@ B.1 Details for Screw Theory
For a revolute joint, the screw axis satisfies $\omega||=1$ and $v=-\omega\times q$, where $q$ is an arbitrary point on the screw axis. For a prismatic joint, $\omega=0$ and v=1||v||=1. Given a screw axis $\mathcal{S}$ and a joint angle $\theta$, the motion of an arbitrary rigid body coordinate $T\in\mathrm{SE}(3)$ along the screw axis can be expressed using the matrix exponential:
@@
where $T^{\prime}\in\mathrm{SE}(3)$ denotes the transformed rigid body coordinate, and $\mathcal{S}]$ is the $4\times 4$ matrix representation of the screw axis $\mathcal{S}$, defined as
$\mathcal{S}]=\begin{bmatrix}[\omega]&v\\ 0&0\end{bmatrix}\in\mathbb{R}^{4\times 4},\>\>\>[\omega]=\begin{bmatrix}0&-\omega_{3}&\omega_{2}\\ \omega_{3}&0&-\omega_{1}\\ -\omega_{2}&\omega_{1}&0\end{bmatrix}\in\mathbb{R}^{3\times 3},$ (12)
@@
For a revolute joint, $G(\theta)$ has a closed-form expression using the fact that $\omega]^{3}=[\omega]$:
@@ B.2 Implementation Details for ScrewSplat
Initialization of Core Components. For the part-aware Gaussian primitives, we initially spawn 10,000 primitives. For each primitive $\mathcal{H}_{i}=(T_{i},s_{i},\sigma_{i},c_{i},m_{i})$, we generally adopt the same initialization scheme as used in the original Gaussian Splatting method for the variables $(T_{i},s_{i},\sigma_{i},c_{i})$. It is worth noting that the position $\mu_{i}$ of the Gaussian primitive from the pose $T_{i}=[R_{i},\mu_{i}]$ is sampled from a uniform distribution: in the simulation environment, it is sampled from $1,1]^{3}$, , and in real-world experiments, it is sampled from $0.7,0.7]^{3}$. We found that initializing the positions within the camera’s range results in better performance. The part probability $m_{i}\in\Delta^{n_{s}}$ is initialized as a uniform distribution (i.e., an element-wise constant vector).
For the screw primitives, we initially spawn eight revolute joint axes and eight prismatic joint axes (i.e., a total of 16 joint axes). For each screw primitive $\mathcal{A}_{j}=(\mathcal{S}_{j},\gamma_{j})$, we first sample a six-dimensional real vector from a uniform distribution over $0.5,0.5]^{3}$ and then normalize it to satisfy the constraints for both revolute and prismatic joints. Specifically, let the sampled six-dimensional vector be $[x,q]$, where $x\in\mathbb{R}^{3}$ and $q\in\mathbb{R}^{3}$. For a revolute joint $\omega,v]$, $\omega$ is set to x/xx/||x|| and $v$ is set to $\omega\times q$. For a prismatic joint $\omega,v]$, $\omega$ is set to the zero vector, and $v$ is set to q/qq/||q||. The screw confidence $\gamma_{j}$ is initialized to 0.9.
@@ B.3 Details for Articulated Object Manipulation
Robot Trajectory Planning for Real-world Articulated Object Manipulation. Given the current joint angle $\theta_{c}\in\mathbb{R}$ and target joint angle $\theta_{t}\in\mathbb{R}$, along with a screw primitive $\mathcal{A}_{j}=(\mathcal{S}_{j},\gamma_{j})$, we propose a simple robot trajectory planning approach for real-world articulated object manipulation. The trajectory planning consists of two main stages: first, planning the robot gripper’s tip trajectory, and second, planning the gripper’s $\text{SE}(3)$ trajectory based on the tip trajectory.
@@
After designing the tip trajectory, we then plan the gripper’s $\text{SE}(3)$ trajectory, ensuring that the gripper’s tip follows the tip trajectory while maintaining a fixed orientation. Typically, we set the orientation so that the robot gripper faces toward the ground. Once the $\text{SE}(3)$ trajectory is obtained, we solve the inverse kinematics to compute the final robot joint angle trajectory for articulated object manipulation.
@@ C.1 Additional Details for Recognition Experiments
Evaluation Metrics. We again note that we adopt three types of metrics, including geometry, motion, and appearance, for articulated object recognition experiment. For geometry metric, we first reconstruct meshes from the recognized models. Specifically, for PARIS, PARIS*, and DTA, we use the marching cubes method [49]. For ScrewSpawn and ScrewSplat, we render depth images from multiple designated camera views and fuse them using the Truncated Signed Distance Function (TSDF). The final mesh is then extracted using the marching cubes method from the TSDF. From each reconstructed mesh, we uniformly sample 2,048 points to obtain a point cloud and compute the bi-directional Chamfer-$l_{2}$ distances as described in Section 5.1.